Circular No. 8641

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005lr IN ESO 492-G2 Further to IAUC 8634, M. Baek and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images. SN 2005lr is located at α = 7^{h}11^{m}39^{s}.03, δ = −26^{o}42′20′′.2 (equinox 2000.0), which is 18′′.8 west and 2′′.1 south of the nucleus of the galaxy ESO 492-G2. Available magnitudes: 2003 Oct. 29.45 UT, [19.5; 2005 Oct. 25.40, [19.0; Nov. 5.41, [18.5; Dec. 4.41, 18.5; 5.39, 18.5.

V2361 CYGNI C. C. Venturini, R. J. Rudy, D. K. Lynch, and S. Mazuk, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; R. B. Perry, Langley Research Center, NASA; and B. Walp, Lick Observatory, report 0.47- to 2.5-μm spectroscopy of V2361 Cyg (cf. IAUC 8483, 8487) with the Lick 3.0-m telescope (+ VNIRIS) at Nov. 13.212 UT. The near-infrared excess has diminished, which suggests a lowering of the dust temperature. The object is not as heavily extincted as was previously observed, probably a result of dissipation of a dust shell formed after outburst. The lower-excitation coronal lines of [Si VI] and [Ca VIII] are present. He I 1.0830-μm is dominant and He II lines are present. [N I] is identified, and the unidentified novae lines are weakly present. The optical spectrum is faint, and only Hα, [O III], and [N II] are prominent — although other weaker features are present.

NOVA IN THE LARGE MAGELLANIC CLOUD 2005 Additional CCD V magnitudes (± 0.06) from W. Liller (cf. IAUC

8635): Dec. 4.071 UT, V = 11.96; 5.077, 12.03; 6.067, 11.88.

COMET P/2005 W3 (KOWALSKI)

Improved elliptical orbital elements from MPEC 2005-X24:

T = 2005 Aug. 23.0791 TT  ω = 199.^{o}2400
e = 0.530569  Ω = 211.5687 }2000.0
q = 3.008609 AU  i = 16.7793

a = 6.409057 AU n^{o} = 0.0607453 P = 16.22 years

2005 December 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green